Solve
sqrt(x^2+7x-2) -3=x

Answer by edjones(8007)

sqrt(x^2+7x-2) -3=x
sqrt(x^2+7x-2) =x+3
x^2+7x-2=x^2+6x+9 square each side.
x=11
.
Ed